# Building Guardrails for Large Language Models

### v1

@@ 1 Introduction
Recent times have witnessed a notable increase in the utilization of Large Language Models (LLMs) like ChatGPT, attributed to their extensive and general capabilities (OpenAI, 2023). However, the rapid deployment and integration of LLMs have raised significant concerns regarding their risks including, but not limited to, ethical use, data biases, privacy and robustness (Huang et al., 2023b). In societal contexts, worries also include the potential misuse by malicious actors for activities such as spreading misinformation or aiding criminal activities, as indicated in studies by Kreps et al. (2022); Goldstein et al. (2023); Kang et al. (2023). In the scientific context, LLMs can be used in professional contexts, where there are dedicated ethical considerations and risks in scientific research (Birhane et al., 2023).
@@
Like safety-critical software, a systematic process to cover the development cycle (ranging from specification, to design, implementation, integration, verification, validation, and production release) is required to carefully build the guardrails, as indicated in industrial standards such as ISO-26262 and DO-178B/C. The goal of this paper is to review the state-of-the-art (Section 2), present technical challenges on implementing individual requirements (Section 3), and then discuss several issues regarding the systematic design of a guardrail for a specific application context (Section 4).
@@ 2 Existing Implementation Solutions
$Nvidia\ NeMo$, described in (Rebedea et al., 2023), functions as an intermediary layer that enhances the control and safety of LLMs. It employs Colang, an executable programme language designed by Nvidia (2023), to establish constraints, in order to guide LLMs within set dialogical boundaries. When the customer’s input prompt comes, NeMo embeds the prompt as a vector, and then uses K-nearest neighbor (KNN) method to compare it with the stored vector-based user canonical forms, retrieving the embedding vectors that are ‘the most similar’ to the embedded input prompt. After that, Nemo starts the flow execution to generate output from the canonical form. During the flow execution process, the LLMs are used to generate a safe answer if requested by the Colang program. The process is presented in Figure 2. Building on the above customizable workflow, NeMo also includes a set of pre-implemented moderations dedicated to e.g., fact-checking, hallucination prevention in responses, and content moderation. NeMo is also a Type-1 neural-symbolic system, with its effectiveness closely tied to the performance of the KNN method.
@@ 3.4 Hallucinations and Uncertainty
Protection via I/O Engineering Apart from the refining methods, Pinter & Elhadad (2023) found that these methods might pose potential risks when trying to combat LLMs hallucinations. They recommend using retrieval-augmented methods, which seek to add external knowledge acquired from retrieval directly to the LLMs’ prompt (He et al., 2022; Press et al., 2022; Ram et al., 2023). Based on the Chain-of-Thought technology, Dhuliawala et al. (2023) introduced the “Chain-of-Verification” method to effectively reduce the generation of inaccurate information in LLMs. Wang et al. (2023a) then proposed a faithful knowledge distillation method that significantly enhances the credibility and accuracy of LLMs. Zhao et al. (2023) proposed a Verify-and-Edit framework based on GPT-3, which enhances the factual accuracy of predictions in open-domain question-answering tasks. Additionally, Gao et al. (2023) pioneered the “Retrofit Attribution using Research and Revision” system, which improves the outputs by automatically attributing and post-editing generated text to correct inaccuracies.
@@ 4.1 Conflicting Requirements
This section discusses the tension between safety and intelligence, as an example for the conflicting requirements. Conflicting requirements are typical, including e.g., fairness and privacy (Xiang, 2022), privacy and robustness (Song et al., 2019), and robustness and fairness (Bassi et al., 2024).
@@
The integration of guardrails with LLMs may lead to a discernible conservative shift in the generation of responses to open-ended text-generation questions (Röttger et al., 2023). The shift has been witnessed in ChatGPT over time. Chen et al. (2023) documented a notable change in ChatGPT’s performance between March and June 2023. Specifically, when responding to sensitive queries, the model’s character count decreased significantly, plummeting from an excess of 600 characters to approximately 140. Additionally, in the context of opinion-based questions and answers surveys, the model is more inclined to abstain from responding.
@@
Given the brevity and conservativeness of responses generated by ChatGPT, it raises the question: How can exploratory depth be maintained in responses, particularly for open-ended test generation tasks? Furthermore, does the application of guardrails constrain ChatGPT’s capacity to deliver more intuitive responses? On the other hand, Narayanan & Kapoor (2023) critically examined this paper, and emphasized the difference between an LLM’s capabilities and its behavior. Although capabilities typically remain constant, behavior can alter due to fine-tuning, which can be interpreted as the “uncertainty” challenges in LLMs. They suggest that changes in GPT-4’s performance are likely linked more to evaluation data and fine-tuning methods rather than a decline in its fundamental abilities. They also acknowledge that such behavioral drift poses a challenge in developing reliable chatbot products. The adoption of guardrail has also led to the model adopting a more succinct communication approach, thereby offering fewer details or electing for non-response in certain queries. The decision of “to do or not to do” can be a challenging task when designing the guardrail. While the easiest approach is to decline an answer to any sensitive questions, is it the most intelligent one? That is, we need to determine if the application of guardrail always has a positive impact on LLMs that is within our expectation.
@@
Our Perspective Prior research has suggested to incorporate a creativity assessment mechanism into the guardrail development for LLMs. To measure the creativity capability of LLMs, Chakrabarty et al. (2023) employed the Consensual Assessment Technique (Amabile, 1982), a well-regarded approach in creativity evaluation, focusing on several key aspects: fluency, flexibility, originality, and elaboration, which collectively contribute to a comprehensive understanding of the LLMs’ creative output in storytelling. Narayanan & Kapoor (2023) showed that although some LLMs may demonstrate adeptness in specific aspects of creativity, there is a significant gap between their capabilities and human expertise when evaluated comprehensively.
@@ 4.3 Neural-Symbolic Approach for Implementation
Our Perspective First, a principled approach is needed to resolve conflicts in requirements, as suggested in (van Lamsweerde et al., 1998) for requirement engineering, which is based on the combination of logic and decision theory. Second, a guardrail requires the cooperation of symbolic and learning-based methods. For example, we may expect that, the learning agents deal with the frequently-seen cases (where there are plenty of data) to improve the overall performance w.r.t. the above-mentioned requirements, and the symbolic agents take care of the rare cases (where there are few or no data) to improve the performance in dealing with corner cases in an interpretable way. Due to the complex conflict resolution methods, more closely-coupled neural-symbolic methods might be needed to deal with the tension between effective learning and sound reasoning, such as those Type-6 systems (Lamb et al., 2021) that can deal with true symbolic reasoning inside a neural engine, e.g., Pointer Networks (Vinyals et al., 2015).
@@ 4.4 Systems Development Life Cycle (SDLC)
Our Perspective Rigorous verification and testing will be needed (Huang et al., 2023b), which requires a comprehensive set of evaluation methods. For individual requirements, certification with statistical guarantees can be useful, such as the randomized smoothing (Cohen et al., 2019). For the evaluation of multiple, conflicting requirements, a combination of the Pareto front based evaluation methods for multiple requirements (Ngatchou et al., 2005) and the statistical certification for a single requirement is needed. Finally, attention should also be paid to understanding the theoretical limits of the evaluation methods, e.g., randomized smoothing causes a fairness problem (Mohapatra et al., 2021).
@@ 5 Conclusion
This position paper advocates for a systematic approach to building guardrails, beyond the current solutions which only offer the simplest mechanisms to describe rules and connect learning and symbolic components. Guardrails are highly complex due to their role of managing interactions between LLMs with humans. A systematic approach, supported by a multidisciplinary team, can fully consider and manage the complexity and provide assurance to the final product.
@@ Appendix B Demonstration of the current challenges in ChatGPT
Figure 4: Harmful Response

### v2

@@ 1 Introduction
Recent times have witnessed a notable increase in the utilization of Large Language Models (LLMs) like ChatGPT, attributed to their extensive and general capabilities (OpenAI, 2023). However, the rapid deployment and integration of LLMs have raised significant concerns regarding their risks including, but not limited to, ethical use, data biases, privacy and robustness (Huang et al., 2023d). In societal contexts, worries also include the potential misuse by malicious actors for activities such as spreading misinformation or aiding criminal activities, as indicated in studies by Kreps et al. (2022); Goldstein et al. (2023); Kang et al. (2023). In the scientific context, LLMs can be used in professional contexts, where there are dedicated ethical considerations and risks in scientific research (Birhane et al., 2023).
@@
This paper argues that, like safety-critical software, a systematic process to cover the development cycle (ranging from specification, to design, implementation, integration, verification, validation, and production release) is required to carefully build the guardrails, as indicated in industrial standards such as ISO-26262 and DO-178B/C. The goal of this paper is to review the state-of-the-art (Section 2), present technical challenges on implementing individual requirements (Section 3), and then discuss several issues regarding the systematic design of a guardrail for a specific application context (Section 4).
@@ 2 Existing Implementation Solutions
$Nvidia\ NeMo$, described in (Rebedea et al., 2023), functions as an intermediary layer that enhances the control and safety of LLMs. NeMo is designed as a versatile toolkit that facilitates the creation, training, and deployment of state-of-the-art LLMs, including but not limited to GPT. LLMs are extensively used throughout the guardrail process for various tasks across multiple stages. For example, in a conversation scenario, LLM is utilized in the following three phases: (I) Generating user intent, where it refines user intent using provided examples and potential intents, producing deterministic results by setting the temperature to zero. (II) Generating next step: In this phase, Nemo searches the most relevant similar flows, and integrates these similar flows together into an example, which is then fed into the LLM. The output of LLM call is termed as “bot intent”. (III) Generating the bot message, taking the most relevant five bot intents and relevant data chunks as inputs to provide context.
Unlike traditional models that rely on initial layer embeddings, NeMo utilizes similarity functions to capture the most pertinent semantics, employing the “sentence transformers / all-MiniLM-L6-v2” model for this purpose. This model aids in embedding inputs into a dense vector space, enhancing the efficacy of nearest neighbor searches using the Annoy algorithm. Additionally, NeMo employs Colang, an executable programme language designed by Nvidia Nvidia (2023), to establish constraints, in order to guide LLMs within set dialogical boundaries. When the customer’s input prompt comes, NeMo embeds the prompt as a vector, and then uses K-nearest neighbor (KNN) method to compare it with the stored vector-based user canonical forms, retrieving the embedding vectors that are ‘the most similar’ to the embedded input prompt. After that, Nemo starts the flow execution to generate output from the canonical form. During the flow execution process, the LLMs are used to generate a safe answer if requested by the Colang program. The process is presented in Figure 2. Building on the above customizable workflow, NeMo also includes a set of pre-implemented moderations dedicated to e.g., fact-checking, hallucination prevention in responses, and content moderation. NeMo is also a Type-1 neural-symbolic system, with its effectiveness closely tied to the performance of the KNN method.
@@
Overall, in this section we review three existing strategies for implementing guardrails, each with its own set of pros and cons. Subsequent sections will delve into methodologies for constructing guardrail components, tailored to meet specific requirements. Especially, Section 3 provides an overview of the current research landscape on individual requirements, and Section 4 delivers a broader systems-thinking approach to consider multiple requirements altogether.
@@ 3.4 Hallucinations and Uncertainty
Protection via I/O Engineering Apart from the refining methods, Pinter & Elhadad (2023) found that these methods might pose potential risks when trying to combat LLMs hallucinations. They recommend using retrieval-augmented methods, which seek to add external knowledge acquired from retrieval directly to the LLMs’ prompt (He et al., 2022; Press et al., 2022; Ram et al., 2023). Based on the Chain-of-Thought technology, Dhuliawala et al. (2023) introduced the “Chain-of-Verification” method to effectively reduce the generation of inaccurate information in LLMs. Wang et al. (2023b) then proposed a faithful knowledge distillation method that significantly enhances the credibility and accuracy of LLMs. Zhao et al. (2023) proposed a Verify-and-Edit framework based on GPT-3, which enhances the factual accuracy of predictions in open-domain question-answering tasks. Additionally, Gao et al. (2023) pioneered the “Retrofit Attribution using Research and Revision” system, which improves the outputs by automatically attributing and post-editing generated text to correct inaccuracies.
@@ 4.1 Conflicting Requirements
This section discusses the tension between safety and intelligence, as an example for the conflicting requirements. Conflicting requirements are typical, including e.g., fairness and privacy (Xiang, 2022), privacy and robustness (Song et al., 2019), robustness and XAI (Huang et al., 2023c), and robustness and fairness (Bassi et al., 2024). The integration of guardrails with LLMs may lead to a discernible conservative shift in the generation of responses to open-ended text-generation questions (Röttger et al., 2023). The shift has been witnessed in ChatGPT over time. Chen et al. (2023) documented a notable change in ChatGPT’s performance between March and June 2023. Specifically, when responding to sensitive queries, the model’s character count decreased significantly, plummeting from an excess of 600 characters to approximately 140. Additionally, in the context of opinion-based questions and answers surveys, the model is more inclined to abstain from responding.
@@
Given the brevity and conservativeness of responses generated by ChatGPT, it raises the question: How can exploratory depth be maintained in responses, particularly for open-ended test generation tasks? Furthermore, does the application of guardrails constrain ChatGPT’s capacity to deliver more intuitive responses? On the other hand, Narayanan & Kapoor (2023) critically examined this paper, and emphasized the difference between an LLM’s capabilities and its behavior. In psychological studies (Michie et al., 2011), behaviour is believed to be determined by not only capability (refer to knowledge, skills, etc) but also opportunity for external factors and motivation for internal processes. In the context of LLMs, the opportunity includes social norms and cultural practices that need to be taken care of by the guardrails. Although capabilities typically remain constant, behavior can alter due to fine-tuning, which can be interpreted as the “uncertainty” challenges in LLMs. They suggest that changes in GPT-4’s performance are likely linked more to evaluation data and fine-tuning methods rather than a decline in its fundamental abilities. They also acknowledge that such behavioral drift poses a challenge in developing reliable chatbot products. The adoption of guardrail has also led to the model adopting a more succinct communication approach, thereby offering fewer details or electing for non-response in certain queries. The decision of “to do or not to do” can be a challenging task when designing the guardrail. While the easiest approach is to decline an answer to any sensitive questions, is it the most intelligent one? That is, we need to determine if the application of guardrail always has a positive impact on LLMs that is within our expectation.
@@
Our Perspective For the safety and intelligence tension, prior research has suggested to incorporate a creativity assessment mechanism into the guardrail development for LLMs. To measure the creativity capability of LLMs, Chakrabarty et al. (2023) employed the Consensual Assessment Technique (Amabile, 1982), a well-regarded approach in creativity evaluation, focusing on several key aspects: fluency, flexibility, originality, and elaboration, which collectively contribute to a comprehensive understanding of the LLMs’ creative output in storytelling. Narayanan & Kapoor (2023) showed that although some LLMs may demonstrate adeptness in specific aspects of creativity, there is a significant gap between their capabilities and human expertise when evaluated comprehensively. We also need to assess which requirements are critical and which can be adjusted or compromised for different tasks and contexts. While these conflicts may not be entirely resolvable, particularly within a general framework applicable across various contexts, more targeted approaches in specific scenarios might offer better chance of conflict resolution. Such approaches demand ongoing research to develop concrete principles, methods, and standards that a multidisciplinary team can implement and adhere to. Guardrails, while effective in particular situations, are not a universal solution capable of addressing all potential conflicts. Instead, they should be designed to manage specific, well-defined scenarios.
@@ 4.3 Neural-Symbolic Approach for Implementation
Our Perspective First, a principled approach is needed to resolve conflicts in requirements, as suggested in (van Lamsweerde et al., 1998) for requirement engineering, which is based on the combination of logic and decision theory. Second, a guardrail requires the cooperation of symbolic and learning-based methods. For example, we may expect that, the learning agents deal with the frequently-seen cases (where there are plenty of data) to improve the overall performance w.r.t. the above-mentioned requirements, and the symbolic agents take care of the rare cases (where there are few or no data) to improve the performance in dealing with corner cases in an interpretable way. In general, before we can confirm, and reliably evaluate, the cognitive ability of learning agents, the symbolic agents can embed human-like cognition (e.g., the analogical connections between concepts in similar abstract contexts) through structures such as knowledge graphs. Not only can they improve the guardrails’ capability, but they also enable the end users with more explainability, which is important due to the guardrails’ responsibility in providing safety and trust to AI. Due to the complex conflict resolution methods, more closely-coupled neural-symbolic methods might be needed to deal with the tension between effective learning and sound reasoning, such as those Type-6 systems (Lamb et al., 2021) that can deal with true symbolic reasoning inside a neural engine, e.g., Pointer Networks (Vinyals et al., 2015).
@@ 4.4 Systems Development Life Cycle (SDLC)
Our Perspective Rigorous verification and testing will be needed (Huang et al., 2023d), which requires a comprehensive set of evaluation methods. For individual requirements, certification with statistical guarantees can be useful, such as the randomized smoothing (Cohen et al., 2019) and global robustness (Dong et al., 2023). For the evaluation of multiple, conflicting requirements, a combination of the Pareto front based evaluation methods for multiple requirements (Ngatchou et al., 2005) and the statistical certification for a single requirement is needed. The Pareto front, a concept from the field of multi-objective optimization, represents a set of non-dominated solutions, where no solution is better than others across all objectives that are considered. Some efforts have been taken, e.g., (Huang et al., 2023c) adapts an evolutionary algorithm to find Pareto front for robustness and XAI. Statistical certification involves using statistical methods to ensure that a single requirement meets a specified standard with a certain level of confidence. It is typically applied when there is uncertainty in the measurements or when the requirement is subject to variability. Combining these techniques can find the trade-offs, provide confidence in the viability of solutions with respect to individual requirements, and support more informed and adaptive decision-making processes. Attention should also be paid to understanding the theoretical limits of the evaluation methods. For example, it is known that different verification methods will provide different levels of guarantees on their results, with (”davidad” Dalrymple et al., 2024) defining 11 levels (0-10), e.g., the commonly applied attacks are only at level-1, some testing methods (Sun et al., 2019; Wicker et al., 2018) are at level-5 or level-6, and methods based on sampling with global optimisation guarantees or statistical guarantees such as (Cohen et al., 2019; Dong et al., 2023; Ruan et al., 2018; Wang et al., 2023a) are at between level-7 and level-9. Last but not least, safety argument (Zhao et al., 2020; Dong et al., 2023) will be needed to not only structure the reasoning and evidence collection but also ensure the communication with the stakeholders.
@@ 5 Conclusion
This paper advocates for a systematic approach to building guardrails, beyond the current solutions which only offer the simplest mechanisms to describe rules and connect learning and symbolic components. Guardrails are highly complex due to their role of managing interactions between LLMs with humans. A systematic approach, supported by a multidisciplinary team, can fully consider and manage the complexity and provide assurance to the final product.
Acknowledgements
This project has received funding from The Alan Turing Institute under grant agreement No ARC-001 and the U.K. EPSRC through End-to-End Conceptual Guarding of Neural Architectures [EP/T026995/1].
@@ Appendix B Demonstration of the current challenges in ChatGPT
Figure 4: Harmful Response on ChatGPT 3.5